OpenStudy (anonymous):

Line AB contains points A (−2, 6) and B (4, 5). The slope of line AB is 1. negative 1 over 6 2.1 over 6 3.6

OpenStudy (anonymous):

2.

OpenStudy (anonymous):

like hell

OpenStudy (candyme):

The slope of a line is\[\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\] So plug in your y values in the top and your x values in the bottom

OpenStudy (anonymous):

\[\huge \frac{6-5}{4-(-2)}\] will ifnd it

OpenStudy (firekat97):

remember \[slope = \frac{ y2 - y1 }{ x2 - x1 }\]

OpenStudy (anonymous):

i got 5-6 over 4-(-2) so what do i do next?

OpenStudy (candyme):

Subtract them 5-6=1 4-(-2)=4+2=6 So the slope is....

OpenStudy (anonymous):

6?

OpenStudy (candyme):

Don't forget the top

OpenStudy (anonymous):

oh.. 1 over 6?

OpenStudy (candyme):

there you go

OpenStudy (anonymous):

yay! thank you!!

OpenStudy (candyme):

Oh actually I mistyped that myself. Should be -1/6

OpenStudy (candyme):

Sorry about that